# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

### v1

@@ 4.2 Evaluation
We evaluate on a diverse set of four knowledge-intensive NLP tasks. PopQA (Mallen et al., 2023) is a free-formed long-tail open-domain question answering dataset. Following Asai et al. (2024), we use the subset of 1,399 questions that aims at testing knowledge of long-tail entities. TriviaQA (Joshi et al., 2017) is an established open-domain question answering dataset that features relatively complex and diverse questions. We use the same test split and retrieval setup as in Asai et al. (2024). PubHealth (Zhang et al., 2023) is a fact-checking dataset focusing on checking claims in the public health domain. Finally, ARC Challenge (Bhakthavatsalam et al., 2021) is a multiple-choice question answering dataset featuring grade-school level science questions. Following common practice, we perform lexical postprocessing of the model’s output and report accuracy for PubHealth and ARC and substring matching for PopQA and TriviaQA.
@@
We compare SR-RAG with baselines that cover various training and inference strategies. (1) First, using the LLM before fine-tuning, we compare with either always retrieving or always verbalizing with GenRead. (2) As illustrated in Figure 2, the major baseline we compare with is the state-of-the-art prior selective retrieval pipeline, combining the advantage of He et al. (2021), Asai et al. (2024), and Wu et al. (2024a). Specifically, the likelihoods of the LLM generating the answer with and without retrieval are used to create the knowledge source selection label. Then, we fine-tune the LLM for knowledge source selection (among $S_{e}$ and $\phi$) and generate the answer with optional retrieval. At inference, we apply a uniform threshold of 0.2 on the likelihood of the retrieval token for selective retrieval. (3) Always retrieving with the fine-tuned SR-RAG LLM.
@@ 5.1 Knowledge Verbalization Alters Knowledge Source Preference
Table 1: Main evaluation results on four tasks. The best results are boldfaced. Across three LLMs as the backbone models, SR-RAG consistently outperforms selective RAG and always retrieving while consuming a much lower retrieval budget.
To motivate our approach, we first demonstrate that knowledge verbalization has a substantial impact on identifying when an LLM requires retrieval. Figure 3 presents a pilot study on Llama-2-7B-Chat using four datasets from SR-RAG’s training mixture. For each instance, we compute the likelihood of the LLM generating the correct answer with no context (blue), with the most helpful GenRead verbalization ($c_{i+}$, orange), and with the most helpful retrieved passage ($c_{e+}$, red). Notably, GenRead changes the preferred knowledge source in 16% of Natural Questions examples and over 30% in the other datasets. This suggests that prior selective retrieval methods, which omit verbalization, may significantly underestimate the LLM’s capabilities—reinforcing the need to incorporate knowledge verbalization for accurate source preference labeling.
5.2 SR-RAG Improves Accuracy and Reduces Retrieval
Table 1 shows the end-to-end generation performance on three LLMs, demonstrating the advantage of SR-RAG over always retrieving, selective retrieval, and other baselines. Relative to always retrieving with the pretrained model, SR-RAG improves accuracy by 19.3% (Llama-2-7B-Chat), 13.4% (Phi-3.5-Mini-Instruct), and 5.9% (Qwen2.5-7B-Instruct). Against the strongest selective retrieval baseline, SR-RAG yields 8.5%/2.1%/4.7% higher accuracy. Interestingly, despite saving 8–15% retrievals, the selective retrieval baseline achieves accuracy nearly identical to always retrieving. This suggests that selective RAG alone cannot reliably identify retrieval-unnecessary queries. In contrast, SR-RAG reduces 20–40% of low-value retrievals while maintaining or improving performance due to accurate knowledge source selection and high-quality verbalization.
Remarkably, with a uniform inference datastore and threshold, SR-RAG dynamically adapts its retrieval behavior based on dataset characteristics. For a dataset that emphasizes long-tail knowledge like PopQA, SR-RAG tends to retrieve external knowledge most of the time. On the other hand, for PubHealth and ARC where the model’s knowledge may suffice for a number of questions, SR-RAG relies on internal knowledge more confidently, resulting in much better performance compared to always retrieving. Qualitative examples illustrating SR-RAG’s behavior can be found in Figure 9 and Figure 10 in the appendix.
5.3 SR-RAG Makes Accurate Source Selection Decisions
@@
AUROC (Retrieval $>$ Verbalization)
@@
Table 2: Source selection accuracy measured on Llama-2-7B-Chat. SR-RAG achieves the best averaged performance in both settings.
Can SR-RAG accurately choose the right knowledge source? Table 2 compares SR-RAG with Self-RAG (Asai et al., 2024) and an SR-RAG variant without kNN-based policy. We evaluate accuracy using two criteria: (1) abstaining from retrieval when it does not harm performance (top) and (2) retrieving only when it provides strictly better context (bottom). While all methods perform well under the first criterion, SR-RAG achieves the highest overall accuracy. Under the second, stricter definition, SR-RAG demonstrates substantially better discrimination, outperforming Self-RAG by 14.7% in average AUROC. Removing kNN-based policy (w/o. kNN) leads to noticeable drops in both accuracy and AUROC, highlighting the benefit of adapting to fine-tuning-induced shifts in LLM ability.
5.4 System Efficiency
@@
We assess SR-RAG’s end-to-end latency using Llama-2-7B-Chat under batched inference [^5] Detailed latency modeling formulation in Section B.2. In our setting, latency is linearly correlated with retrieval frequency under small batch sizes.. Figure 4 shows the trade-off between accuracy and latency under varying retrieval proportions. As expected, latency improves as fewer instances invoke retrieval. Due to dataset-specific difficulty, the optimal retrieval proportion varies per dataset. However, SR-RAG’s learned source selection policy achieves a near-optimal accuracy-efficiency trade-off across all datasets, without any dataset-specific threshold tuning. This confirms that SR-RAG not only improves performance but does so efficiently.
5.5 Ablation Study
@@
To understand the contribution of each SR-RAG component, we perform ablations on Llama-2-7B-Chat (Table 3). Removing kNN-based source selection (w/o. kNN) results in lower accuracy and higher retrieval rates, confirming that kNN inference helps the model adapt to ability shifts introduced by fine-tuning. In addition, disabling knowledge verbalization during preference labeling (w/o. kv. label) causes the LLM to over-rely on retrieval and reduces performance, showing that verbalization is crucial for effective labeling. Finally, $\mathcal{L}_{stage2}$ is ablated and the $\mathcal{L}_{stage1}$ loss is kept for stage 2 training. The result shows a significant increase in retrieval proportion as the model’s knowledge verbalization ability degrades. These findings highlight that all three components are vital to SR-RAG’s success. In Appendix C, we include more detailed analyses on the <EOQ> representations, other SR-RAG hyperparameters, as well as qualitative examples.
@@ Appendix A List of Notations
In Table 4, we present the major notations and parameters used throughout the paper.
@@ B.1 Training Details
Algorithm 1 presents the full algorithm for constructing training data and labeling knowledge source preferences. GenRead is executed independently on each training data subset. We adopt instance-level notation for clarity. The pipeline naturally scales to additional knowledge sources by applying knowledge collection and likelihood evaluation in parallel across sources.
@@
6:     $\mathcal{C}_{S_{i}}\leftarrow$ $M$.GenRead$(q,n)$
@@
We implement GenRead following the setup in Yu et al. (2023). In the second verbalization round, five clusters with five in-context examples each are used. All datasets except ASQA follow the general prompt shown in Figure 5. For ASQA, we include an additional instruction to handle ambiguity: “If the question is ambiguous, generate multiple documents for each possibility.
@@
Table 5 summarizes the training and validation splits, including the proportion of examples where verbalization is preferred over retrieval.
@@ B.2 Latency Formulation
To evaluate the inference efficiency of SR-RAG, we measure the latency in a realistic batched inference setup, where the system handles a batch of $B$ = 10 queries and returns the results for all of them together. For the latency experiments presented in the paper, we decompose the system latency as follows:
  • Source Selection Time ($T_{d}$): The time taken by the knowledge source selector to determine whether to retrieve from external sources or rely on parametric knowledge.
  • Retrieval Latency ($T_{rs}$): The time taken to fetch external knowledge from the database if the model chooses to retrieval from an external knowledge source $s$. In our batched setting, we calculate $T_{rs}$ by performing a batched retrieval for all the $B$ instances that require retrieval and report the per-item latency.
  • Verbalization Latency ($T_{v}$): The time taken for the LLM to verbalize parametric knowledge if itself is selected as the knowledge source.
  • Generation Latency ($T_{g}$): The time required for the LLM to generate the response, conditioned on either retrieved or verbalized knowledge.
@@
$T_{\text{total}}=\begin{cases}T_{d}+T_{v}+T_{g},&\text{if verbalize},\\ T_{d}+T_{rs}+T_{g},&\text{if retrieve from knowledge source $s$}.\end{cases}$
@@
  • We choose the batched setup due to the complexity of the retrieval system. For instance, in our implementation of Wikipedia search, it takes around five to ten seconds per instance to encode the query and retrieve the most relevant context chunks. In the batched setting with queries executed in parallel, the throughput better measures the advantage of SR-RAG.
@@
  • As the nearest neighbor only involves one matrix product for similarity calculation and one top-k operation, $T_{d}$ is generally very small. In fact, we find $T_{d}$ (0.01s) <<much-less-than$<$ < $T_{g}$ (0.1s) $<$ $T_{v}$ (1s). On the other hand, $T_{rs}$ is the major bottleneck of the pipeline. As a result, in an online setting, the system’s efficiency gain directly converges to the percentage of retrievals it is able to avoid.
Appendix C Further Analyses
C.1 Self-Routing RAG: Hyperparameters

### v2

@@ 4.2 Evaluation
We evaluate on a diverse set of four knowledge-intensive NLP tasks. PopQA (Mallen et al., 2023) is a free-formed long-tail open-domain question answering dataset. Following Asai et al. (2024), we use the subset of 1,399 questions that aims to test knowledge of long-tail entities. TriviaQA (Joshi et al., 2017) is an established open-domain question answering dataset that features relatively complex and diverse questions. We use the same test split and retrieval setup as in Asai et al. (2024). PubHealth (Zhang et al., 2023) is a fact-checking dataset focusing on checking claims in the public health domain. Finally, ARC Challenge (Bhakthavatsalam et al., 2021) is a multiple-choice question answering dataset featuring grade-school level science questions. Following common practice, we perform lexical post-process of the model’s output and report accuracy for PubHealth and ARC and substring matching for PopQA and TriviaQA.
@@
We compare SR-RAG with baselines that cover various training and inference strategies. (1) First, using the LLM before fine-tuning, we compare with either always retrieving or always verbalizing with GenRead. (2) As illustrated in Figure˜2, our main baseline is the state-of-the-art prior selective retrieval pipeline, combining the advantage of He et al. (2021), Asai et al. (2024), and Wu et al. (2024a). Specifically, the likelihoods of the LLM generating the answer with and without retrieval are used to create the knowledge source selection label. Then, we fine-tune the LLM for knowledge source selection (among $S_{e}$ and $\phi$) and generate the answer with optional retrieval. At inference, we apply a uniform threshold of 0.2 to the likelihood of the retrieval token for selective retrieval. (3) Always retrieving with the fine-tuned SR-RAG LLM.
@@ 5.1 Knowledge Verbalization Alters Knowledge Source Preference
Table 1: Main evaluation results on four tasks. The best results are boldfaced. Across three backbone LLMs, SR-RAG consistently outperforms selective RAG and always retrieving while using a much lower retrieval budget.
To motivate our approach, we first show that knowledge verbalization substantially impacts on identifying when an LLM requires retrieval. Figure˜3 presents a pilot study on Llama-2-7B-Chat using four datasets from SR-RAG’s training mixture. For each instance, we compute the likelihood of the LLM generating the correct answer with no context (blue), with the most helpful GenRead verbalization ($c_{i+}$, orange), and with the most helpful retrieved passage ($c_{e+}$, red). Notably, GenRead changes the preferred knowledge source in 16% of Natural Questions examples and over 30% in the other datasets. This suggests that prior selective retrieval methods, which omit verbalization, may significantly underestimate the LLM’s capabilities, reinforcing the need to incorporate knowledge verbalization for accurate source preference labeling.
@@
AUROC (Retrieval >> Verbalization)
@@
Table 2: Source selection accuracy measured on Llama-2-7B-Chat. SR-RAG achieves the best averaged performance in both evaluation settings.
@@
5.2 SR-RAG Improves Accuracy with Reduced Retrieval Frequency
Table˜1 shows the end-to-end generation performance on three LLMs, demonstrating the advantage of SR-RAG over always retrieving, selective retrieval, and other baselines. Relative to always retrieving with the pretrained model, SR-RAG improves accuracy by 19.3% (Llama-2-7B-Chat), 13.4% (Phi-3.5-Mini-Instruct), and 5.9% (Qwen2.5-7B-Instruct). Against the strongest selective retrieval baseline, SR-RAG yields 8.5%/2.1%/4.7% higher accuracy. Interestingly, despite saving 8–15% retrievals, the selective retrieval baseline achieves nearly identical accuracy to always retrieving. This suggests that selective RAG alone cannot reliably identify queries that do not require retrieval. In contrast, SR-RAG reduces 20–40% retrievals while maintaining or improving performance thanks to accurate knowledge source selection and high-quality verbalization.
Remarkably, with a uniform inference datastore and threshold, SR-RAG dynamically adapts its retrieval behavior based on dataset characteristics. For a dataset that emphasizes long-tail knowledge like PopQA, SR-RAG tends to retrieve external knowledge most of the time. On the other hand, for PubHealth and ARC where the model’s knowledge may suffice for a number of questions, SR-RAG relies on internal knowledge more confidently, resulting in much better performance compared to always retrieving. Qualitative examples illustrating SR-RAG’s behavior can be found in Figure˜7 and Figure˜8 in the appendix.
5.3 SR-RAG Makes Accurate Source Selection Decisions
Can SR-RAG accurately choose the correct knowledge source? Table˜2 compares SR-RAG with Self-RAG (Asai et al., 2024) and an SR-RAG variant without kNN-based policy. We evaluate accuracy using two criteria: (1) abstaining from retrieval when it does not harm performance (top) and (2) retrieving only when it provides strictly better context (bottom). While all methods perform well under the first criterion, SR-RAG achieves the highest overall accuracy. Under the second, stricter definition, SR-RAG demonstrates substantially better performance, outperforming Self-RAG by 14.7% in average AUROC. Removing kNN-based policy (w/o. kNN) leads to large drops in both accuracy and AUROC, highlighting the benefit of adapting to fine-tuning-induced shifts in LLM ability.
5.4 System Efficiency
We assess SR-RAG’s end-to-end latency using Llama-2-7B-Chat under batched inference [^6] Detailed latency modeling formulation in Section B.2. In our setting, latency is linearly correlated with retrieval frequency under small batch sizes.. Figure˜4 shows the trade-off between accuracy and latency under varying retrieval proportions. As expected, latency improves as fewer instances invoke retrieval. Due to dataset-specific difficulty, the optimal retrieval proportion varies per dataset. However, SR-RAG’s learned source selection policy achieves a near-optimal accuracy-efficiency trade-off across all datasets, without any dataset-specific threshold tuning. This confirms that SR-RAG not only improves performance but does so efficiently.
5.5 Ablation Study
To understand the contribution of each component, we perform ablations on Llama-2-7B-Chat (Table˜3). Removing kNN-based source selection (w/o. kNN) results in lower accuracy and higher retrieval rates, confirming that kNN inference helps the model adapt to ability shifts introduced by fine-tuning. Disabling knowledge verbalization during preference labeling (w/o. kv. label) causes the LLM to over-rely on retrieval and reduces performance, showing that verbalization is crucial for effective labeling. Finally, $\mathcal{L}_{stage2}$ is ablated and the $\mathcal{L}_{stage1}$ loss is kept for stage 2 training. The result shows significantly more retrievals as the model’s knowledge verbalization ability degrades. These findings highlight that all components are vital to SR-RAG’s success. In Appendix˜D, we further analyze the <EOQ> representations, other SR-RAG hyperparameters, as well as qualitative examples.
@@ Appendix A List of Notations
In Table˜4, we present the major notations and parameters used throughout the paper.
@@ B.1 Training Details
Algorithm˜1 presents the full algorithm for constructing training data and labeling knowledge source preferences. GenRead is executed independently on each training data subset. We adopt instance-level notation for clarity. The pipeline naturally scales to additional knowledge sources by applying knowledge collection and likelihood evaluation in parallel across sources.
@@
6:   $\mathcal{C}_{S_{i}}\leftarrow$ $M$.GenRead(q,n)(q,n)
@@
We implement GenRead following the setup in Yu et al. (2023). In the second verbalization round, five clusters with five in-context examples each are used. All datasets except ASQA follow the general prompt shown in Figure˜5. For ASQA, we include an additional instruction to handle ambiguity: “If the question is ambiguous, generate multiple documents for each possibility.
@@
Table˜5 summarizes the training and validation splits, including the proportion of examples where self-verbalized knowledge is preferred over retrieved knowledge.
@@ B.2 Latency Formulation
To evaluate the inference efficiency of SR-RAG, we measure the latency in a realistic batched inference setup, where the system handles a batch of $B$ = 10 queries and returns the results for all queries. For the latency experiments presented in the paper, we decompose the system latency as follows:
  • Source Selection Time ($T_{d}$): The time for the knowledge source selector to determine whether to retrieve from external sources or rely on parametric knowledge.
  • Retrieval Latency ($T_{rs}$): The time for fetching external knowledge from the database if the model chooses to retrieval from an external knowledge source $s$. In our batched setting, we calculate $T_{rs}$ by performing a batched retrieval for all the $B$ instances that require retrieval and report the per-item latency.
  • Verbalization Latency ($T_{v}$): The time for the LLM to verbalize parametric knowledge if itself is selected as the knowledge source.
  • Generation Latency ($T_{g}$): The time for the LLM to generate the response, conditioned on either retrieved or verbalized knowledge.
@@
$T_{\text{total}}=\begin{cases}T_{d}+T_{v}+T_{g},&\text{if verbalize},\\ T_{d}+T_{rs}+T_{g},&\text{if retrieve from $s$}.\end{cases}$
@@
  • We choose the batched setup due to the complexity of the retrieval system. For instance, in our implementation of Wikipedia search, it takes around five to ten seconds per instance to encode the query and retrieve the most relevant context chunks. The batched setting amortizes the latency of single retrieval.
@@
  • As the nearest neighbor only involves one matrix product for similarity calculation and one top-k operation, $T_{d}$ is generally very small. In fact, we find $T_{d}$ (0.01s) <<<< $T_{g}$ (0.1s) << $T_{v}$ (1s). On the other hand, $T_{rs}$ is the major bottleneck of the pipeline. As a result, in an online setting, the system’s efficiency gain directly converges to the percentage of retrievals it is able to avoid.
Appendix C SR-RAG: Multi-Source Extensions
Beyond routing between the LLM itself and an external knowledge source, SR-RAG’s formulation is naturally compatible with routing between multiple external knowledge sources as well. In this section, we provide further experiments on routing between three knowledge sources and discuss other potential multi-source applications.
C.1 Experiment: Routing between General and Specialized Sources
We perform experiments on the PubMed corpus (National Center for Biotechnology Information (), NCBI), which contains professional biomedical literature. Compared to Wikipedia and the LLM itself, PubMed has the unique advantage of featuring long-tail specialized biomedical knowledge. Meanwhile, it falls short at providing knowledge for many general factuality questions, posing a challenge to accurately routing the knowledge source.
Based on this new setup, we formulate SR-RAG with three knowledge source tokens: <Self>, <Wiki>, and <Pubmed>. We leverage the MedRAG toolkit (Xiong et al., 2024) to implement retrieval on PubMed. The training and inference strategies of the two-source SR-RAG recipe naturally extends to this formulation and are followed in our experiments here. For training and evaluation, we additionally use PubMedQA (Jin et al., 2019), a widely used biomedical question answering dataset. PubMedQA features short questions extracted from PubMed abstracts and provides associated ground truth yes/no/maybe answers. We sample 10k artifically generated documents for training (combined with the data mixture used in the main text) and 500 examples from the expert labeled subset for evaluation. The same set of LLMs from three families (Llama, Qwen, Phi) are trained and evaluated.
In Table˜6, we present the three-way routing results of SR-RAG and compare with always RAG as well as GenRead with the model before fine-tuning. Our main findings are as follows:
  • Adding PubMed as an additional knowledge source does not harm its effectiveness. The results on datasets other than PubMedQA aligns with the performance of two-source SR-RAG reported in the main text.
  • SR-RAG training improves the model’s ability to verbalize its knowledge on PubMedQA as well while being much more efficient than GenRead, as shown by the better performance of SR-RAG + Always Verbalizing compared to GenRead.
  • SR-RAG intelligently routes the best knowledge source for different workloads. For PubMedQA, PubHealth, and ARC, where PubMed can outperform Wiki by a large margin, SR-RAG frequently self-selects PubMedQA as its preferred knowledge source. On the other hand, on general factuality benchmarks such as PopQA and TriviaQA, SR-RAG selects Wikipedia more frequently. More importantly, SR-RAG achieves this behavior end-to-end without the need for any manual interventions.
  • On PubMedQA, the end-to-end SR-RAG achieves the best performance compared to the baselines. The trend is consistent across three different LLMs.
Overall, this experiments confirms the effectiveness of extending SR-RAG to three sources and suggests a promising outlook for further extension.
Training Inference PubMedQA PopQA TriviaQA PubHealth ARC Average
Llama-2-7B-Chat
No Fine-tuning GenRead 0.544 0.247 0.616 0.515 0.605 0.505
Always RAG, Wiki 0.636 0.529 0.641 0.457 0.546 0.562
Always RAG, PubMed 0.656 0.157 0.470 0.612 0.537 0.486
SR-RAG (3-source) Always RAG, Wiki 0.342 0.587 0.639 0.668 0.639 0.572
Always RAG, PubMed 0.672 0.325 0.539 0.713 0.645 0.642
Always Verbalizing 0.388 0.305 0.572 0.724 0.631 0.579
SR-RAG 0.680 0.583 0.647 0.724 0.655 0.677
(% <Wiki>) 6% 95% 88% 17% 4% 29%
(% <Pubmed>) 91% 4% 7% 70% 86% 64%
(% <Self>) 3% 1% 5% 13% 10% 8%
Phi-3.5-Mini-Instruct
No Fine-tuning GenRead 0.376 0.331 0.567 0.442 0.840 0.511
Always RAG, Wiki 0.460 0.541 0.594 0.549 0.771 0.583
Always RAG, PubMed 0.676 0.143 0.356 0.447 0.726 0.470
SR-RAG (3-source) Always RAG, Wiki 0.486 0.580 0.634 0.676 0.805 0.636
Always RAG, PubMed 0.738 0.302 0.494 0.644 0.813 0.598
Always Verbalizing 0.498 0.327 0.510 0.704 0.819 0.572
SR-RAG 0.736 0.560 0.639 0.716 0.824 0.695
(% <Wiki>) 1% 95% 87% 54% 0% 47%
(% <Pubmed>) 97% 3% 10% 18% 27% 31%
(% <Self>) 2% 2% 3% 28% 73% 22%
Qwen2.5-7B-Instruct
No Fine-tuning GenRead 0.418 0.334 0.626 0.676 0.875 0.586
Always RAG, Wiki 0.612 0.563 0.667 0.446 0.916 0.641
Always RAG, PubMed 0.698 0.129 0.469 0.401 0.821 0.504
SR-RAG (3-source) Always RAG, Wiki 0.358 0.576 0.639 0.637 0.818 0.606
Always RAG, PubMed 0.726 0.307 0.503 0.676 0.824 0.607
Always Verbalizing 0.522 0.303 0.550 0.695 0.828 0.580
SR-RAG 0.730 0.572 0.651 0.681 0.834 0.694
(% <Wiki>) 15% 96% 86% 20% 0% 43%
(% <Pubmed>) 79% 1% 13% 48% 39% 36%
(% <Self>) 6% 3% 1% 32% 61% 21%
Table 6: Three-way routing between retrieving from Wikipedia, retrieving from PubMed, or self-verbalizing knowledge. For each model we report task accuracy and the selection frequencies of the individual sources wiki/PubMed/self under SR-RAG’s LLM+kNN knowledge source routing scheme.
C.2 Discussion: Further Applications
SR-RAG has numerous useful real-world applications and exhausitively experimenting on them is beyond the scope of this paper’s experiments. In this section, we discuss several potential applications that SR-RAG naturally supports.
Cost-Aware Routing
Querying different knowledge sources brings distinct latency cost. SR-RAG’s preference label can be constructed such that the cheapest useful knowledge source is preferred, improving the latency while preventing the performance degradation. In addition, varied levels of compute are required to verbalize the required knowledge for different types of queries. SR-RAG can be used to model this pattern by formulating multiple internal knowledge sources associated with different levels of compute. This philosophy connects with the literature on adaptive reasoning budget (Lou et al., 2025).
Recency-Aware Routing
Many user queries have an inherent temporal dimension (e.g., “today,” “as of Sept. 2025,” fast-moving entities) where stale knowledge degrades correctness. SR-RAG can expose “freshness-tiered” sources and learn a preference label that prioritizes the most up-to-date tier when temporal cues or out-of-vocabulary entities are detected. The router can condition on time expressions, named-entity recency features, and source timestamps. When the model’s self-estimate is uncertain or predates the query’s time scope, it escalates to a recency-guaranteed index.
Privacy-Aware Routing
Queries often mix public facts with sensitive tenant data (e.g., customer tickets, internal codenames, emails). SR-RAG can represent privacy tiers as distinct sources and train preference labels under a “least-privilege first” policy: prefer private indices when identifiers or confidentiality cues are present and fall back to public only when the query is non-sensitive and private stores are irrelevant.
Appendix D Further Analyses
D.1 Self-Routing RAG: Hyperparameters